A meeting of envoys was attended by 4 Koreans and 2 Filipinos. If two envoys were



selected at random one after the other, determine the values of the random variable K



representing the number of Koreans. Construct the probability histogram.

1
Expert's answer
2022-03-17T06:36:54-0400

Solution

There are 4 Koreans and 2 Filipinos.

Let;

K represent the number of Koreans

F represent the number of Filipinos


If two people are chosen at random one after the other there are 3 possible outcomes


  1. Both the selected people are Filipinos and hence "K=0"
  2. One Korean and one Filipino is chosen hence "K=1"
  3. Both of the selected people are Koreans and hence "K=2"


The probability values of K are as follows


"K=0:~P(F,F)"


"=\dfrac26\times\dfrac15=\dfrac{1}{15}"


"K=1: ~P(K,F) or P(K,F)"


"=(\dfrac46\times\dfrac25)+(\dfrac26\times\dfrac45)=\dfrac{8}{15}"



"K=2:~P(K,K)"


"=\dfrac46\times\dfrac35=\dfrac25"


The probability histogram